. A 74.2 kg teenager slides down a frictionless slide that has an angle of inclination of 38°. What is the magnitude of the acceleration of the teenager down the slide?


1
Expert's answer
2021-02-22T16:03:18-0500
"a=g\sin{38}\\a=9.8\sin{38}=6.03\frac{m}{s^2}"